Grade 12th passVectors

If |A×B| = (A.B) , then what is the value of |A+B|.

ravi
8 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear Ravi
 

|A×B|/A.B=√3

=> ABsin x / ABcos x= √3 (Considering the angle between the two vectors as x)

=> tanx =√3

=> x= π/3

=> |A+B|=√(A^2 + B^2 +2AB cos x)

=> |A+B|=√(A^2 + B^2 +2AB* 1/2) [Since cos π÷3 = 1/2]

=> |A+B|=√(A^2 + B^2 + AB)

 

Regards

Arun (askIITians forum expert)